3.6
Solve and check the following Linear equation
2=5(x-4)-(x-6)
What is the solution?
The solution set is { }

Answer by Theo(13342)

2 = 5(x-4) - (x-6)

remove parentheses to get 2 = 5x - 20 - x + 6

combine like terms to get 2 = 4x - 14.

add 14 to both sides of the equaton to get 2 + 14 = 4x

simplify to get 16 = 4x.

divide both sides of the equation by 4 to get x = 4.

replace x with 4 in your original equation to get 2 = 5(x-4) - (x-6) becomes 2 = 5(4-4) - (4-6) which becomes 2 = 5*0 - (-2) which becomes 2 = 0 + 2 which becomes 2 = 2, confirming the solution is correct.